OpenStudy (anonymous):

Solve using the square root method. x2 = 64

OpenStudy (anonymous):

\[x^{2} = 64\]\[\sqrt{x^{2}} = \sqrt{64}\]\[x = ±8\]

OpenStudy (anonymous):

\[x^2 = 64\implies x= \pm8\]

OpenStudy (anonymous):

thank you guys! :)